OpenStudy (anonymous):

Evaluate lim x->+infinity of [(3x^3+2x-5)/( sqrt of x^2+x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow +\infty} (3x ^{3}+2x - 5) / (\sqrt{x ^{2}+x})\]

OpenStudy (anonymous):

infinity with your eyeballs

OpenStudy (anonymous):

numerator is a polynomial of degree 3 and while the denominator is not a polynomial, \[\sqrt{x^2}=x\] so the denominator behaves like a polynomial of degree 1, and since 3 is bigger than 1 you get infinity

OpenStudy (anonymous):

inside the radical it is x^2 +x not only x and i also need to show work